OpenStudy (anonymous):

What is the simplified form of 5sqrtx times itself 4 times??? WILL GIVE MEDAL!!

TheSmartOne (thesmartone):

Is it\[(5\sqrt{x})^4\]

TheSmartOne (thesmartone):

\[5\sqrt{x}\times 5\sqrt{x}\times 5\sqrt{x}\times 5\sqrt{x}= (5\sqrt{x})^4\]

OpenStudy (anonymous):

Um, I think so. But instead of the exponant there, they just have (5sqrtx) listed four times with the multiplication symbol in the middle of each one.

TheSmartOne (thesmartone):

Like what I just posted

OpenStudy (anonymous):

Yes, exactly!

TheSmartOne (thesmartone):

Do you understand what that 4 at the top right hand corner means.

OpenStudy (anonymous):

no, i don't really understand this problem at all. this was on a pre-test, so i've never learned this stuff before. i told my teacher that i have no idea how to anser these kinds of problems, but he wants me to try and answer them anyway. :/

TheSmartOne (thesmartone):

That means whatever is in the parenthesis multiplied to itself four times.

OpenStudy (anonymous):

okay, so how do you multiply a problem like this?

TheSmartOne (thesmartone):

You need to know that \[5\sqrt{x}=5(x)^\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

okay...

TheSmartOne (thesmartone):

So it will become \[(5(x)^\frac{ 1 }{ 2 })^4\]

TheSmartOne (thesmartone):

\[5^4\times x^(\frac{ 1 }{ 2}\times 4)\]

TheSmartOne (thesmartone):

\[625\times x^2\]

TheSmartOne (thesmartone):

Because 1/2 times 4 = 2

TheSmartOne (thesmartone):

and 5^4 =5x5x5x5 =625

OpenStudy (anonymous):

does that mean 625x^2 is the answer? or the simplified form? cause this question also asks for the simplified form, and 625x^2 is not one of the choices listed...

TheSmartOne (thesmartone):

Simplified form would become \[(25x)^2\]

OpenStudy (anonymous):

that's not one of the options either. the ones they have listed are 4^5sqrtx, 5 sqrt4x, x^5/4 and x^4/5.

TheSmartOne (thesmartone):

You could make it \[(5^2x)^2\]

TheSmartOne (thesmartone):

Or just \[5^4x^2\]

OpenStudy (anonymous):

hmm, well then maybe there's an error on the test? because your answers aren't matching what they have. I appreciate the help though. :)

TheSmartOne (thesmartone):

Ask your teacher. He/She can give you back points for this question..

OpenStudy (anonymous):

okay, thank you!